OpenStudy (anonymous):

Solve the area formula for a triangle, A = 1/2 bh, for h

OpenStudy (anonymous):

what no numbers ?

OpenStudy (anonymous):

Nope this is all they gave me

OpenStudy (texaschic101):

A = 1/2 bh for h Multiply both sides by 2 (2)A = (2)1/2 bh 2A = bh Now divide both sides by b (2A)/b = bh/b (2A)/b = h

OpenStudy (anonymous):

Right thats the formula you reverse or should i text reverse the problem

OpenStudy (anonymous):

Thanks!

OpenStudy (texaschic101):

no problem...glad I could help :)